OpenStudy (tiffany_rhodes):

solve y' = (x+1)/(y-1)

OpenStudy (tiffany_rhodes):

I started with integrating both sides:

zepdrix (zepdrix):

Start by separating variables, yes? :)

OpenStudy (tiffany_rhodes):

\[\int\limits_{?}^{?}(y-1)*dy/dx=\int\limits_{?}^{?}(x+1)\]

OpenStudy (tiffany_rhodes):

Yeah, I multiplied (y-1) over to the dy/dx side

zepdrix (zepdrix):

And then "multiply" dx to the other side, and integrate, Looks good so far.

OpenStudy (tiffany_rhodes):

Alright. So \[\int\limits_{?}^{?}(y-1)*dy = \int\limits_{?}^{?}(x+1)*dx\]

OpenStudy (tiffany_rhodes):

and then integrate?

zepdrix (zepdrix):

Yup :) Understand how to integrate each term? :o Just power rule for the most part.

OpenStudy (tiffany_rhodes):

yes :) thanks! I got